### v1

# Knowledge-Guided Prompt Learning for Request Quality Assurance in Public Code Review

Abstract.
Public Code Review (PCR) is an assistant to the internal code review of the development team, in the form of a public Software Question Answering (SQA) community, to help developers access high-quality and efficient review services. Current methods on PCR mainly focus on the reviewer’s perspective, including finding a capable reviewer, predicting comment quality, and recommending/generating review comments. However, it is not well studied that how to satisfy the review necessity requests posted by developers which can increase their visibility, which in turn acts as a prerequisite for better review responses. To this end, we propose a Knowledge-guided Prompt learning for Public Code Review (KP-PCR) to achieve developer-based code review request quality assurance (i.e., predicting request necessity and recommending tags subtask). Specifically, we reformulate the two subtasks via 1) text prompt tuning which converts both of them into a Masked Language Model (MLM) by constructing prompt templates using hard prompt; 2) knowledge and code prefix tuning which introduces external knowledge by soft prompt, and uses data flow diagrams to characterize code snippets. Finally, both of the request necessity prediction and tag recommendation subtasks output predicted results through an answer engineering module. In addition, we further analysis the time complexity of our KP-PCR that has lightweight prefix based the operation of introducing knowledge. Experimental results on the PCR dataset for the period 2011-2023 demonstrate that our KP-PCR outperforms baselines by 8.3%-28.8% in the request necessity prediction and by 0.1%-29.5% in the tag recommendation. The code implementation is released at https://github.com/WUT-IDEA/KP-PCR.
public code review, review necessity prediction, tag recommendation, prompt learning, external knowledge
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 8ccs: Software and its engineering Software post-development issues
1. Introduction
Code review can be conducted either within a team or by public developers. As an application form of traditional code review (Shan et al., 2022; Chen et al., 2022b; Li et al., 2023a; Rahman et al., 2017; Pandya and Tiwari, 2022), code review within a team, serves as an essential quality assurance mechanism in mature and successful commercial and Open Source Software (OSS) projects (Bosu and Carver, 2013). However, due to its labor-intensive nature, code review within the team gradually increases the time and resources spent on review activities as projects become more complex and extensive (Rigby and Bird, 2013; Rigby et al., 2014). Such cost requirements slow down the popularity of team code review (Johnson, 1998). Consequently, modern code review has evolved from a formal and rigorously regulated process (Sadowski et al., 2018; Fagan, 2011) to a less stringent practice (Baum et al., 2017a, 2016; Rigby and Bird, 2013).
Public Code Review (PCR), as an assistant to internal code review within teams, takes the form of Software Question Answering (SQA) community to assist developers in exploring high-quality and efficient review services (Hong et al., 2022). Developers have widely adopted it as a new option for code review. Research has found that the code review process not only helps to improve the overall quality of code and software systems (Ackerman et al., 1989; Baum et al., 2017b), identify defects (Baum and Schneider, 2016), but also promotes the exchange of professional and development technical knowledge (Sadowski et al., 2018). As an emerging form of service, in PCR developers publish review requests within a community for all community members to view, then the reviewers complete the review service through comments. In the PCR process implemented by the SQA community, request quality assurance for code review service mainly depends on developers. As developers are required and encouraged to submit PCR requests in the community, the necessity of these code review requests is evaluated by public via request necessity prediction (Ochodek et al., 2022; Chen et al., 2022a; Hong et al., 2022) subtask after they are posted. Meanwhile, developers always need to choose technical terms as review tags. These tags can be used to match requests with suitable reviewers via tag recommendation (Li et al., 2020; Nie et al., 2020; Xu et al., 2021; He et al., 2022; Li et al., 2023b) subtask.
Existing methods on PCR are mostly based on pre-trained language models. However, the language models lack the learning of professional knowledge vocabulary in the field of PCR. Early work regarded code snippets as non-linguistic characters that would affect the model’s effectiveness, for example, the TagDC method proposed by Li et al. (Li et al., 2020) and the PROFIT model proposed by Nie et al. (Nie et al., 2020). These works have treated code snippets as noise and directly discarded code snippets. Recent studies have discovered semantic information embedded in code snippets, but these studies mostly employ the same processing methods as they do with text (Xu et al., 2021; He et al., 2022). Overall, existing work has not fully utilized the knowledge embedded in code snippets.
To this end, we propose Knowledge-guided Prompt learning for Public Code Review (KP-PCR). Specifically, we first design a task-descriptive prompt template that restructures the request necessity prediction subtask and the tag recommendation subtask into a generative task through a knowledge-guided prompt template, and input the request body (including tags and descriptions) into the prompt construction module of a language model. We reformulate both subtasks via 1) text prompt tuning which converts two subtasks into Masked Language Model (MLM) by constructing prompt templates using hard prompt; and 2) knowledge and code prefix tuning which introduces external knowledge by soft prompt, and uses data flow diagrams to characterize code snippets. Finally, the request necessity prediction and recommendation tags output predicted results through an answer engineering module. In addition, we delve deeper into the time complexity analysis of our KP-PCR with lightweight prefix based operations to introducing knowledge, which reveals that our KP-PCR can improve task accuracy performance without compromising overall efficiency. Experimental results demonstrate that our KP-PCR outperforms baselines in terms of accuracy-related metrics across both subtasks. Furthermore, we explore the impact of prompt learning construction on the results, and use a case study to compare our KP-PCR with GPT-3.5 and GPT-4.0 considering tag recommendation.
The main contributions of this study can be summarized as follows:
  • From the perspective of developers, to satisfy the request quality assurance, a unified method, called KP-PCR, is proposed for the request necessity prediction subtask and the tag recommendation subtask.
  • We introduce the external knowledge from Wikipedia to enhance the representation of code snippets in PCR to address the semantic understanding of domain knowledge in developer PCR requests.
  • Experimental results on the PCR dataset for the time span 2011-2023 demonstrate that our KP-PCR outperforms the baselines of two request quality assurance subtasks.
This paper extends the preliminary version of our study, as presented in (Chen et al., 2024). In comparison, this extended version (1) introduces external knowledge from Wikipedia to generate prefix vectors as soft prompt for the representation of code snippets, (2) designs the knowledge-guided prompt learning on the request necessity prediction subtask and the tag recommendation subtask to show that the knowledge guide can further improve the ability of our KP-PCR method in dealing with the code snippets, (3) comprehensively evaluates the effects of our task-descriptive prompt template to explore the influence of external knowledge on unifiedly modelling the above two subtasks, (4) analyzes the time complexity to prove that our KP-PCR can improve the accuracy of request necessity prediction and tag recommendation while maintaining the overall efficiency of our KP-PCR method, and (5) conducts a case study to investigate the performance of our prompt template on large language models. Additionally, we have examined the challenges of how to design hard and soft prompt templates and delved further into the originally explored unified modelling in the initial study. More details on the extension points compared to our preliminary study in (Chen et al., 2024) can be found in Section 7.5.
The remainder of this paper is structured as follows. Section 2 presents a motivating example, and Section 3 introduces the task addressed in this paper. Section 4 details the proposed method. Section 5 provides the experimental setup, and Section 6 discusses the experiment results and presents a case study to explore the utilization of external knowledge by our KP-PCR. Section 7 describes the related work. Section 8 concludes this work with future directions.
2. Motivating Example
Public Code Review (PCR) is developed in the Software Question Answering (SQA) community. Researchers are increasingly exploring ways to achieve high-quality and efficient review services (Hong et al., 2022; Ackerman et al., 1989; Baum et al., 2017b; Baum and Schneider, 2016; Sadowski et al., 2018). A review request consists of several parts, as shown in Fig. 1, including title, request description (including text and code snippets), quality, and tags.
Figure 1. An Example of a Public Code Review Request.
[…]
the request body, tags, and request contributions
Since PCR differs from team internal code review in service form, the quality of the PCR process mainly depends on the developers who write and submit the review requests rather than the reviewers (Xu et al., 2021; Li et al., 2023b). Existing research is highly effective for review quality, but there is still a room for improvement. Due to the lack of pre-learning of domain knowledge by pre-trained language models, existing methods inadequately learn the domain knowledge implicit in request description.
One major characteristic of the request body is the abundance of code snippets, which is particularly significant for PCR. Specifically, for the task of assessing the necessity in code reviews, it is required by the rules of PCR websites that the code snippets in the requests should constitute runnable and correct code; otherwise, the code snippets fail to meet the review standards. As for the task of tag recommendation, even large language models that currently excel in various aspects still showed pay attention on professional knowledge tags. These factors affecting task performance are closely related to the professional knowledge contained in both text and code snippet semantics.
Figure 2. An Example of the Subject Matter Involving Professional Knowledge in a Request.
[…]
The code snippet can be clearly identified to the java language information.
Code snippets are beneficial for understanding and distinguishing the requests of public code reviews. Taking a specific review request as an example, as shown in Fig. 2, code snippets can clearly identify Java language information. By selecting processing and representation methods that are consistent with the semantic structure of the code snippets, the variable flow expression of the code snippets can be better understood, revealing the information embedded in the logic of the Java programming language. However, the professional knowledge and logical expressions contained in these code snippets may not fully pre-learned by text-based language models.
The text of review requests may contain some vocabulary that is not present in the language model. When tokenizing the text, the language model may split these out-of-vocabulary words into subwords or letters, which limit the model’s performance in completing public code review tasks. Table 1 illustrates an example of applying specialized knowledge in the context of the tag recommendation subtask. As shown in this example, a review request contains the abbreviation “OOP” to describe the content, which is not covered in the language model. However, its meaning is semantically consistent with the label “Object oriented”. Since the language model has not learned “OOP”, it may struggle to associate the abbreviation “OOP” with the label “Object oriented”. In contrast, in knowledge repositories such as Wikipedia, the definition content would link these two terms. Thus, this has motivated us to carry out this study that aims to bridge this gap.
Table 1. Example of Professional Knowledge Guidance
Data Tag Wikipidia
An online store, being converted from procedural to OOPI just start to learn OOP, and it’s far more interesting than procedural style.I have a complete working online store written in procedural style… object oriented Object-oriented programming (OOP) is a programming paradigm based on the concept of objects, which …
3. Task Description
In the PCR process conducted by the SQA community, the quality assurance of requests primarily relies on developers. Developers are required to submit necessary PCR requests to the community, and the necessity of these requests is assessed publicly via the request necessity prediction subtask after submission. At the same time, developers need to select several technical terms as review tags, which can be used to match requests with appropriate reviewers through the tag recommendation subtask.
In this section, we will first describe a public code review process, and then outline how these request quality assurance subtasks (request necessity prediction and tag recommendation) serve the public code review process. This process consists of five core steps (as shown in Fig. 3), starting from the submission of changes or optimization requests by the code developer in Step ①, to the submission of reviews by reviewers in Step ⑤.
3.1. Public Code Review Process
First, the public code review process will be introduced through an example of public code review. As shown in Fig. 3, a review process can be abstracted into 5 steps, labeled as ①-⑤.
Figure 3. An Example of a Public Code Review Process.
[…]
a review process can be abstracted into 5 step
Code review typically involves two roles: the developer who submits the code request and the reviewer who provides suggestions and comments. In Step ①, the developer writes and submits a review request. In Step ②, practitioners in the community review and evaluate the requests to allocate necessary feedback. Subsequently, in Step ③, developers are asked to select several relevant technical terms to accompany their request. These tags are then used in Step ④ to match the request with appropriate reviewers, who provide their feedback on the code in Step ⑤.
In a review request, there are often many changes and optimizations needed in different parts of the code. Reviewers need to spend a significant amount of time reviewing all the changes. However, it is often the case that most changes are minor and do not require comments or suggestions. Li et al. (Li et al., 2022) initially considered the task of assessing code change quality in the internal team code review process, which involves predicting whether a code change is of high quality and ready to be accepted in the review process. This work enhances the efficiency and quality of the review process by pre-evaluating before the request is reviewed. Therefore transitioning to the public code review process, there is a task need for predicting request necessity. Request necessity prediction learns from the existing conventions of public code review and community feedback to determine the standards for submitting review requests, thereby ensuring review quality and reducing the time cost of unnecessary requests.
The task of tag recommendation is common in the software engineering community, typically used for labeling requests and enhancing visibility. In public code review, reviewers browse request content and conduct evaluations based on these tags. Therefore, appropriate tag recommendations contribute to improving review efficiency.
Overall, our task logic will focus on two processes initiated by developers, aiming to optimize the services of public code review. This includes integrating the task of predicting request necessity with Step ②, to provide necessary feedback to practitioners in advance. This can guide them to optimize request expressions and improve request visibility while reducing the time spent by reviewers in Step ⑤. Additionally, the tag recommendation task is associated with Step ④, allowing requests to be correctly matched with appropriate reviewers based on the usage of tags in the PCR community.
3.2. Request Quality Assurance Task Definition
The effectiveness of the PCR process depends on the interaction between developers and reviewers. However, this interaction can be time-consuming due to the need to identify issues, find suitable reviewers, etc. While current methods focus on reducing the time cost for reviewers, they overlook the impact of developers on the review process. For the PCR process, code review requests that comply with review standards are crucial for high-quality review responses and a conducive community atmosphere. To address this issue, this study proposes a quality assurance service for public code review, mainly relying on the request necessity prediction subtask and the tag recommendation subtask.
When considering a request in public code review, it is assumed that there exists a corpus of public code review requests (denoted as $R$), a set of tag labels (denoted by $L$), and a set of quality labels for all PCR requests (denoted by $Q$). Typically, given a public code review request ${req}_{i}\in{R}$, the request consists of a title ${title}_{i}\in{T}$, and a description ${des}_{i}\in{D}$, where the description ${des}_{i}$ includes both text ${tex}_{i}$ and code snippets ${code}_{i}$.
(1) The subtask of request necessity prediction. The objective of the request necessity prediction is to obtain the function quality that maps $r$ to the necessity label ${q}\in{Q}$. A public code review request ${req}_{i}$ corresponds to a necessity label ${q}_{r}$. Therefore, the necessity prediction subtask can be regarded as a classification problem. The goal is to determine an optimal function $f$ that maps the necessity label $q$ to a label similar to the actual necessity label ${q}_{r}$.
(2) The subtask of tag recommendation. The objective of the tag recommendation subtask is to obtain a function that maps ${req}_{i}$ to a set of labels $l=\left\{l_{1},l_{2},\ldots,l_{m}\right\}\subset L$ that are most relevant to the request $r$. A review request $req_{i}$ corresponds to several labels, where $l_{r}=\left\{l_{r1},l_{r2},\ldots,l_{rk}\right\}\subset l$, $k$ represents the number of true labels of the review request. Therefore, the tag recommendation subtask can be viewed as a multi-label classification problem. The goal is to identify an optimal function $f_{tag_{r}ec}$ that maps the label subset $l$ to the true label set $l_{r}$ as closely as possible.
(3) Unified task definition. We design a descriptive prompt template for the two subtasks, summarized as: $T\left(\cdot\right)=``x\left[MASK\right]$” (where $x$ represents a text string). The input sequence $r$ of the request is passed through this descriptive prompt and refactored into $r_{prompt}=T\left(r\right)$. The request necessity prediction and tag recommendation subtasks are made by filling in the position of $\left[MASK\right]$. Our unified task is how to improve the proximity of the predicted filler words against the truth label. We denote the total number of training examples as $N$, the total number of available labels as $M$, and the number of labels in the training data as $l$. The number of $MASK$ is determined by the maximum length of the labels in the two subtasks.
4. KP-PCR Method
In this section, we propose a knowledge-guided prompt learning for PCR called KP-PCR to address the issue of insufficient learning of professional knowledge in public code review.
4.1. Framework Overview
Fig. 4 illustrates the overall framework of KP-PCR which has two modules, knowledge-guided prompt learning module and answer engineering module.
The knowledge-guided prompt learning module consists of text prompt, code prompt and knowledge-guided prompt. First, the request text is input into prompt learning module. The text prompt reconstructs the request text into a generative task, and the code prompt refactors the code snippets in text into the form of data flow graph. These processes are described in more detail in Section 4.2. The request is then matched with the external knowledge base from Wikipedia to get the knowledge guidance, which is then fed into the model in the form of prefix vectors. Further, the obtained prompt template is input into the Masked Language Model (MLM) to get the predicted word list. Then the answer engineering module maps the predicted words to the label list to get the final labels.
Figure 4. The Framework of KP-PCR
[…]
Includes three parts of prompt tuning architecture: text prompt tuning, code prefix tuning, and unified prompt tuning.
4.2. Unified Prompt Template Design
4.2.1. Text Prompt
This subsection elaborates on the process of restructuring subtasks through task descriptive prompt templates, which involves the following two steps:
(1) Task Descriptive Prompt Template
First, a descriptive prompt template $T\left(\cdot\right)$ as a hard prompt is applied to map the input request body $req_{I}=\left\{title_{i},\ des_{i}\right\}$ to prompts. The prompt template retains the original tokens from $req_{i}$ and includes at least one [MASK] for PLM to fill in label words for both subtasks. Specifically, this study modifies the descriptive prompt text based on the subtask objectives to reconstruct the data for different tasks.
Then, by using the descriptive prompt: “The requested label is [MASK][MASK][MASK]. Is this request necessary for review? [MASK]”, the tasks of necessity prediction and tag recommendation are restructured within the same framework.
Finally, with a guiding prompt “Request: ”, the model is directed to explicitly recognize the request content, and concatenate the title $title$, description $des$ (including text content $text$ and code snippet content $code$), and the template $T\left(\cdot\right)$ together. After concatenation, we obtain a text input $req_{i}$ for the model, as shown in Equation 1:
(1) $Input_{i}=\left\{T\left(\cdot\right),title_{i},des_{i}\right\}$
(2) Request Text Encoding
$\left[CLS\right]$ and $\left[SEP\right]$ are special tokens used in BERT-based pre-trained language models. $\left[CLS\right]$ is used to mark the beginning of a sentence, and the representation vector obtained from the encoder can be used for subsequent downstream tasks. $\left[SEP\right]$ is used to mark the end of a sentence or to separate two sentences. The input sequences are fed into a tokenizer which identifies each input with $\left[CLS\right]$ and $\left[SEP\right]$.
Next, as shown in Equation 2, the complete input sequence obtained earlier is fed into the model’s embedding layer for encoding.
(2) $Input_{i}=\left\{\left[CLS\right],T\left(\cdot\right),title_{i},des_{i},\left[ESP\right]\right\}$
The embedding layer of the encoder model initializes each token according to the tokenizer’s tokenization results. Based on pre-learned vector representations, a representation vector for the input sequence is obtained.
4.2.2. Code Prompt
This section mainly introduces the processing of refactoring code snippets into data flow graph based vector. That is used as part of soft prompt.
Code snippets inherently possess a graph structure that provides valuable semantic information for code comprehension. To better capture this information, Guo et al. (Guo et al., 2021) proposed using data flow graphs to analyze code and represent its semantic information. In previous research on code representation, the parsing of code typically involved using Abstract Syntax Trees (ASTs). In contrast, this graph-based data flow structure effectively avoids the unnecessary depth hierarchy that ASTs may introduce.
Figure 5. A Code Snippet Processing Example
[…]
the language of the code snippet is identified.the abstract syntax tree is further parsed into the form of a data flow graph
As shown in Fig. 5, the processing of refactoring code snippets into a data flow graph based vector involves two steps. Firstly, the language of the code snippet is identified, as each programming language has different coding styles and syntax. Therefore, it is necessary to parse the code snippet for each language, extract important nodes and variable information in the form of an abstract syntax tree (AST) after obtaining the language category. Secondly, the abstract syntax tree is further parsed into the form of a data flow graph where nodes are used to record the variable nodes of the code snippet, and edges represent the flow of information between nodes. For example, assigning the value of $b$ to $a$ would be parsed into two nodes, $a$ and $b$, with an edge containing the value of $b$ representing the flow of data.
4.2.3. Knowledge-guided Prompt
In order to optimize the code representation and better guide the knowledge embedded in the request text and code snippets, we use the method of continuous prompting (soft prompt) (Liu et al., 2023b) to introduce external knowledge. Specifically, we first extract the professional term from the request using string matching, then initialize the $P_{prefix}$ vector prefix by concatenating the knowledge interpretation representation, and concatenate the representation vector of the data flow graph of the code snippet after this prefix vector. Finally, the whole vector is represented as the model code snippet instead of the original code snippet. Further details on these extensions can be found in Section 4.3.
4.3. Knowledge-Guided Prompt Learning Module
4.3.1. Knowledge-Guided Prompt Learning
Soft prompting is one of the prompting paradigms, which operates in the embedding space of the model and adjusts the prompt parameters for downstream tasks. Currently, this method has demonstrated superior performance on other non-natural language data, such as image encoding and visual question answering (Liu et al., 2023a; Tsimpoukelli et al., 2021). This indicates that soft prompts are well-suited for learning and optimizing information with such graph structures, owing to the ability of vector embeddings in the embedding space to better adapt to non-natural language structures.
Since the code snippets and text in public code review requests contain a large amount of domain-specific knowledge, which is not fully learned by traditional language models, this study uses the labels of professional terms as a benchmark. According to the vocabulary of pre-trained language models, words not in the vocabulary are first selected, and their professional knowledge definitions are obtained from Wikipedia. These words are then saved as a separate external knowledge file. When processing individual request bodies, since tokenization may split words not present in the language model into roots and suffixes, words are split based on spaces using string manipulation. Furthermore, the split words are subjected to fuzzy casing, tense, and part-of-speech transformations. Next, the processed text string is used to extract several domain-specific knowledge words contained in each request body through string matching. Finally, the definitions of domain-specific knowledge words are concatenated and their representations are obtained, which is then fed into the model in the form of prefix vectors to assist in learning the knowledge in the request body.
Here, taking an input request $req_{i}$ as an example, first, based on the previous data processing, we can obtain two parts of a PCR request:
1. Natural language text $text_{i}$, composed of the title and description (text part) of the request;
2. Code snippet $code_{i}$, representing the code snippet extracted from the description.
Algorithm 1 Knowledge-Guided Prompt Learning
1:$req=\left(title,\ text,\ code\right)$
2:$r_{\widetilde{prompt}}$
3:$Graph_{c}\leftarrow Graph\left(code\right)$, $Graph\left(code\right)$ represents the data flow graph of the code snippet;
4:$h_{Graph_{c}}\leftarrow PLM\left(Graph_{c}\right)$, $h_{Graph_{c}}$ represents the code hidden vector obtained through the encoding model;
5:$knowledge=req\cap\text{wiki}$;
6:$P_{prefix}=PLM\left(knowledge\right),\ PLM\left(knowledge\right)$ represents the representation vector obtained through the encoding model;
7:$r_{\widetilde{prompt}}=\left\{T\left(\cdot\right),\ \widetilde{title},\ \widetilde{text},\ P_{prefix},\ h_{Graph_{c}}\right\}$;
8:return $r_{\widetilde{prompt}}$
The method for processing the code snippet and constructing the prefix vector is shown in Algorithm 1. In Step ①, a code snippet $code_{i}$ is parsed into a data flow graph to represent the dependency relationship among code variables (as shown on the right side of Fig. 5). Then, in Step ②, the variable relationship graph is input into a pre-trained code encoding model for code tasks to obtain the hidden vector of the code snippet. In this work, the GraphCodeBERT (Guo et al., 2021) model is used as the encoding model. In Step ③, professional knowledge terms are extracted from the request using string matching. Step ④ defines a prefix vector $P_{prefix}$, then initializes the prefix vector $P_{prefix}$ with the representation vector of knowledge processed by the encoding model. Finally, the entire vector replaces the original code snippet as the model’s code representation.
Here, $P_{idx}$ is used to represent the sequence of prefix indices, and $\sigma$ is used to represent the non-linear output obtained from the hidden layers of the code pre-trained language model. To address the issue of low performance and high variance caused by the random initialization of $P_{prefix}$, we initialize the prefix vector $P_{prefix}$ with the text of external knowledge and then use an immutable data flow graph based on semantic relationships to represent the meaning of individual code snippets. This allows us to guide the model to learn professional knowledge from the request text through external knowledge and to globally and locally understand the meaning of each code snippet through the form of prefix vectors.
In order to measure the efficiency of Algorithm 1, we analyze its time complexity. To this end, we consider the computational overhead in each step. In Step ①, the size of a code snippet determines the time complexity to convert code into a data flow diagram. If there are $n$ variables or code fragments in the code snippet, generating a dependency diagram requires checking the dependencies between each pair of variables, and consequently the time complexity is $O\left(n^{2}\right)$. Then, in Step ②, the time complexity of GraphCodeBERT generally depends on the architecture of its pre-trained model and the input graph size of a data flow diagram. Overall, the complexity of Step 2 can be seen as $O\left(V+E\right)$, in which $V$ is the number of nodes in the graph, and $E$ is the number of edges. The complexity of this step is also $O\left(n^{2}\right)$, because the encoding process is proportional to the size of the graph. Step ③ uses string matching whose time complexity is $O\left(s\right)$, in which $s$ is the length of the request text. The time complexity of Step ④ depends on the size of the encoding model and the length of the knowledge representation vector. The process of initializing a vector can be regarded as a constant time operation, and the complexity is $O\left(d\right)$, in which $d$ is the length of the vector. Therefore, the overall time complexity of the algorithm can be expressed as $O\left(n^{2}+s+d\right)$. Since the complexity of language model representation is much higher than that of string matching, i.e., $O\left(n^{2}\right)\gg O\left(s\right)$ and $O\left(n^{2}\right)\gg O\left(d\right)$, and the time complexity can also be expressed as $O\left(n^{2}\right)$.
Compared to our preliminary study (Chen et al., 2024) where the time complexity is also $O\left(n^{2}\right)$, it can be found that according to the experimental results in Section 6, our KP-PCR method can improve the task accuracy of request necessity prediction and tag recommendation.
4.3.2. Loss Function
Fig. 6 formalizes the construction of task inputs through an example request, showcasing the joint construction of task description prompt templates (hard prompt) and prefix vectors (soft prompt). It is divided into trainable vector parts and frozen vector parts that are not involved in training.
Figure 6. Trainable and Frozen Parameters
[…]
concatenate the obtained text segment vector and code snippet vector into the input prompt of the Masked Language Model (MLM).
First, the obtained text segment vector and code snippet vector are concatenated into the input prompt of the Masked Language Model (MLM). Within the entire input $\widetilde{c}$ is an immutable vector, while the rest of the vectors are updated based on the gradient optimization target. As shown in Fig. 6, the vector $\widetilde{c}=h_{Graphc}$ is encoded by the hidden layer of the pre-trained code language model is frozen and not updated, whereas the prefix is initialized using the code snippet text. Finally, the vector $h_{PREFIX_{i}}$ can be trained. The dimensions of the code prefix tuning initialization trainable matrix $P_{\sigma}$ (parameterized by $\sigma$) are $|Pidx|\times\text{dim}\left(h_{Xprefixi}\right)$ to store the prefix parameters.
Since multiple [MASK] tokens may appear in the descriptive prompt template, all masked positions are considered for prediction. The final learning objective of our KP-PCR framework is to consider multiple [MASK] tokens in a descriptive prompt template, using the sum of losses from all [MASK] positions to optimize the model. Specifically, the final learning objective of KP-PCR is to maximize Equation 3:
(3) $\log p_{\phi}\left(Lable_{pre}|Req\right)=\prod\limits_{j=1}^{n}{p_{\phi}\left(\left[MASK\right]_{j}=\varphi_{j}\left(label_{r}\right)|T\left(req\right)\right)},$
where $\varphi$ is the predicted probability distribution, and $\varphi_{j}\left(label_{r}\right)$ represents the mapping from the $j$-th $h_{MASK}$ token to the requested label word $label$ in the input sequence. $\phi$ denotes the parameters of the language model that are updatable. $n$ is the number of masked positions in the descriptive prompt template $T\left(req\right)$.
4.4. Answer Engineering Module
After training the model with Equation 3, the final labels need to be obtained, so the answer engineering module maps the predicted words to the label list. The answer engineering module maps the $TOPK$ token list $pre_{list_{i}}$ to the predicted labels $f_{ans}\left(pre_{list_{i}}\right)$ through a mapping function $f_{ans}$. The answer engineering function $f_{ans}$ maps the predicted words to the final label set $Label$. Finally, the mapped answer $f_{ans}\left(pre_{list_{i}}\right)$ is the final top-k outputs $v_{i}$. The mapping function of the answer engineering first compares the predicted words with the label word list to find the complete label by predicting partial label words. Secondly, for data where the predicted words do not exist in the entire label word list, the mapping is completed using the minimum edit distance algorithm. Edit distance, also known as Levenshtein distance, refers to the minimum number of editing operations required to transform one string into another. For example, for the strings “if” and “iff”, you can achieve the transformation by inserting or deleting a “f”.
In general, the smaller the edit distance between two strings, the more similar they are. If two strings are equal, their edit distance is 0 (no operations are needed). In this study, edit distance is used to evaluate the distance between predicted words and true words. If a predicted word is not in the list of true word labels, then by calculating the edit distance between the predicted word and all true words, we can find the true word with the minimum edit distance to the predicted word, thus correcting the predicted word.
5. Experimental Setup
5.1. Research Questions
Our research goal is to provide quality assurance services for PCR from the perspective of developers. Therefore, our method is evaluated from the following three questions.
RQ1: Can our KP-PCR effectively deal with the request necessity prediction subtask and the tag recommendation subtask in a unified modelling?
Our method and SOTA models are applied to examine necessity prediction and tag recommendation subtasks respectively, and their performance on the two subtasks is evaluated.
RQ2: Is our knowledge-guided prompt learning effective for PCR?
Different prompt templates are explored to know the influence of propmt learning on model performance under different designs.
RQ3: How does our unified prompt template perform compared large language models?
We randomly extract 20 request bodies and their corresponding task tags from the existing public code review data. GPT-3.5 and GPT-4.0 are required to select tags according to the task definition in the absence of external resources and with a few references. We explore how GPT-3.5 and GPT-4.0 (1) perform under different prompt descriptions and (2) perform common code review quality assurance tasks by changing the prompt description of the task.
5.2. Dataset and Metrics
5.2.1. Dataset Description
To evaluate the effectiveness of our KP-PCR, a dataset on public code review from Stack Exchange is selected, a reputable well-known software information platform. The selection of the dataset follows the criteria of being open, authoritative, and current. Stack Exchange is a large network of software question and answer (Q&A) websites covering various domains, such as programming, Linux & Unix, and code review. Each website within this network focuses on a specific topic, and the questions, answers, and users are subject to reputation-based recognition within their respective communities.
In the field of software engineering, there are other software engineering information platforms such as Stack Overflow and WordPress. As we are dedicated to providing intelligent service methods for the code review process, the most popular public code review website on Stack Exchange, Code Review [^1] https://codereview.stackexchange.com/, is used as the data source for this study. The dataset from Code Review used for this study is the latest and largest original data version available from 2011 to 2023222https://archive.org/download/stackexchange/codereview.stackexchange.com.7z. It records the main content and quality label of PCR requests in separate XML files.
Fig. 7 introduces the topics of the Code Review website, which is positioned as a peer code review conducted in the form of SQA (Software Quality Assurance). It aims to facilitate a quick and convenient code review process to accelerate the exchange of professional knowledge among peer developers. Any developer can initiate a review request in the community, and peers will assess the necessity of the review based on their expertise, presenting it in the form of contribution scores on each request.
Figure 7. Code Review Website Introduction.
[…]
which is positioned as a peer code review conducted in the form of SQA
5.2.2. Data Preprocessing
To align with the model input requirements of the baseline methods in Section 5.3.2, we have divided the overall data preprocessing into four steps as detailed below.
(1) XML File Parsing: There are two common parsing methods for XML files: 1. DOM parsing which involves converting an XML document into a DOM tree, allowing access to XML document information through a hierarchical object model, describing the information in a node tree format. This parsing method offers flexible access but may suffer from long processing times and memory overflow issues due to machine performance factors. 2. SAX parsing, unlike DOM parsing, operates in a sequential access mode for reading and writing XML files. When using this method to parse XML files, multiple event handling functions need to be written for various data in the file, but it consumes less memory. Due to hardware constraints, this study adopts SAX parsing, resulting in an HTML formatted file.
(2) Data Cleaning and Code Snippet Extraction: Since the XML file contains noise such as URLs, garbled characters, HTML tags, tabs, and line breaks, this step designs regular expressions for the HTML file obtained in Step (1) to remove noise and irrelevant characters. Specifically, the BeautifulSoup library is used to identify HTML tags and separate the text, code snippets, and tags in the request objects. For example, code snippets are collected using the HTML element ¡code¿…¡/code¿.
(3) Tag Filtering and Cleaning: The dataset contains a large number of low-frequency tags. The main reasons for the low frequency of tags are that they do not conform to developers’ usage habits, have weak relevance to the topic, or have low discussion relevance. Previous studies have considered a tag as rare if its frequency is below a predefined threshold $\theta$ (Xu et al., 2021). Intuitively, if a tag appears infrequently in such a large dataset, it is not widely recognized by developers. To mitigate the impact on the quality of training and testing data, these rare tags are removed. Consistent with previous research, this study sets the threshold $\theta$ for rare tags to 50, and all requests and tags that are rare or contain only rare tags are removed. The statistical information of the cleaned dataset is shown in Table 2. Finally, a dataset containing 76,161 requests and 424 tags is obtained.
(4) Tag Processing and Vocabulary Addition: Unlike other topic information stations, the tags in the Code Review involve specialized terms. In addition to common language tags such as “java” and “python,” there are also specific version-related tags such as “c++13” and “python3.x”. The prompt learning method needs to learn and predict based on the words in the vocabulary. Since the model’s tokenization automatically divides words based on symbols, tags are designed as expressions in natural language and added to the model’s vocabulary by writing function rules. This preprocessing of tags is not necessary for baseline methods that do not require the use of a vocabulary for prediction.
Following previous research, we have utilized the common ten-fold cross-validation method to split the dataset. Specifically, we initially randomly have divided the dataset into 10 equally sized subsets. One of these subsets is selected as the test set. The remaining data samples are split into training and validation sets in an 8:2 ratio. This process is repeated during model training to test the model’s effectiveness, and the results is averaged to assess model performance. In summary, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. Detailed statistical information is provided in Table 2.
Table 2. Public Code Review Dataset Statistics
Training Set Validation Set Test Set
All Data 61688 6855 7616
Quality Necessity Task
Necessary 26632 2959 3288
Unnecessary 35056 3896 4328
Tag Recommendation Task
Average Number of Tags / Request Object 2.95 2.89 2.94
Average Number of Words / Tag 2.89 2.89 2.89
5.2.3. Evaluation Metrics
For the tag recommendation subtask, previous recommendation tasks in the SQA community (Li et al., 2023b) have used Precision@k, Recall@k and F1@k to evaluate performance. To maintain consistency, this study adopts the same evaluation metrics. Consistent with the review necessity subtask and the comment quality prediction subtask in the PCR domain (Ochodek et al., 2022), Accuracy and F1 are selected as metrics. The following will provide a detailed introduction to these evaluation metrics.
(1) $Top-k$ Precision $Precision@k$: This metric is commonly used in recommendation-related methods, measuring the proportion of true labels among the top $k$ labels recommended by a model. For a single request, the calculation of its $Top-k$ Accuracy is shown in Equation 4 which computes $Precision@k_{i}$ by calculating the intersection of predicted labels $vk_{i}$ with the corresponding true labels $Ttag_{i}$ for each request object. Since a public code review dataset contains multiple requests, when calculating accuracy metrics for the entire dataset, the average accuracy results for all requests need to be taken into account. The specific calculation method is shown in Equation 5 which represents the average accuracy of all request objects, calculated by first computing $Precision@k_{i}$ for each request object, and then averaging them to represent $Precision@k$.
(4) $Precision@k_{i}=\frac{\left|Ttag_{i}\cap vk_{i}\right|}{k}$
(5) $Presison@k=\frac{\sum\limits_{i=1}^{\left|R\right|}{Precision@k_{i}}}{\left|R\right|}$
(2) $Top-k$ Recall $Recall@k$: This metric is used to evaluate the model’s ability to predict the complete set of true labels. Specifically, it measures the proportion of true labels that are contained in the predicted labels. The $Top-k$ Recall for a single request object, $Recall@k_{i}$, is calculated as shown in Equation 6. In this work, due to the true label length range being 3-5, and following the metrics settings of recommendation-related methods, $k$ is set to 3, 5, and 10 in this study. Similarly, since there are multiple request objects in the public code review dataset, the performance calculation of Recall metric $Recall@k$ for the entire dataset is shown in Equation 7. The average recall rate of all request objects is calculated to represent the model’s performance on this metric.
(6) $Recall@k_{i}=\left\{\begin{array}[]{l}\left|\frac{Ttag_{i}\cap vk_{i}}{k}\right|,if\left|Tag_{i}\right|<k,\\ \left|\frac{Ttag_{i}\cap vk_{i}}{Ttag_{i}}\right|,otherwise.\\ \end{array}\right.$
(7) $Recall@k=\frac{\sum\limits_{i=1}^{\left|R\right|}{Recall@K_{i}}}{\left|R\right|}$
(3) $Top-k$ Macro-Average $F1@k$: This metric is commonly used to evaluate a model’s overall performance on precision and recall by calculating the weighted average of these two metrics. It provides a balanced view of a model’s performance by considering both precision and recall. As described in Equation 8, the calculation method for the macro-average $F1@k_{i}$ of a single request object is similar to the previous two metrics, where $k$ is set to 3, 5, and 10 in this study. Similarly, for the overall macro-average metric $F1@k$ of the public code review dataset, its calculation is shown in Equation 9, derived from the average $F1$ value of all request objects.
(8) $F1@k_{i}=2\times\frac{Precision@k_{i}\times Recall@k_{i}}{Precision@k_{i}+Recall@k_{i}}$
(9) $F1@k=\frac{\sum\limits_{i=1}^{\left|R\right|}{F1@k_{i}}}{\left|R\right|}$
5.3. Implementation Details
5.3.1. Implementation
In practice, the KP-PCR framework can be applied to various pre-trained language models, such as BERT (Beller et al., 2014), RoBERTa (Hong et al., 2022), ALBERT (Thongtanunam et al., 2022), GPT (Hong et al., 2022), etc. However, due to hardware limitations, this study uses the base version of BERT as the encoding model (BERT-base-uncased), and GraphCodeBERT as the code snippet encoder. In the experiments, the learning rate of the KP-PCR is set to 1e-5, and the number of epochs is set to 6.
This study implements KP-PCR based on the PyTorch framework, and all experimental parameters are listed in Table 3. The maximum length of the text is set to 300, the length of the prefix vector is set to 50, and the code representation part of the prefix vector uses a code snippet with a length of 150 to generate the code snippet representation, which is then used for model training and learning with a length of 100. The code snippet text length that is not optimized for model training is 150.
Table 3. Experimental Parameters
Name Parameter Name Parameter
Max lenth 300 Hidden size 768
Prefix lenth 50 Hidden layer 12
Code lenth 150 Weight decay 0
Batch size 4 Learning rate 1e-5
epoch 6 Optimizer Adamw
5.3.2. Baselines
For overall comparisons, a series of state-of-the-art approaches are chosen as baselines for the tag recommendation subtask and the request necessityprediction subtask, respectively.
(1) The request necessity prediction subtask, CommentBERT (Ochodek et al., 2022), MetaTransformer (Chen et al., 2022a) and Commentfinder (Hong et al., 2022) are baseline methods, which have the same data input structure and task definition as review necessity prediction subtask.
(2) The tag recommendation subtask, TagDC (Li et al., 2020), PROFIT (Nie et al., 2020), Post2Vec (Xu et al., 2021), PTM4Tag (He et al., 2022) and CDR4Tag (Li et al., 2023b) are compared for recommending tags to the SQA community.
Our previous study is also listed for comparison. The fine-tuning parameters provided by the baseline methods are used, and their optimal values are adopted in experimenting baselines. For baseline methods which source code was not provided, this study reproduced the model design based on Pytorch framework. The related experiments in this study are conducted on four NVIDIA Titan XP GPUs. All hardware and software environments are shown in Table 4.
Table 4. Experimental Environment and Version Information
Name Version Information
GPU NVIDIA Titan XP*4
CPU IntelR XeonR E5-2650 v4 @2.20GHz
Memory Capacity 64G*4
CUDA Version 11.0
Python Version 3.5
Pytorch Version 1.7
6. Results
6.1. Request Necessity Prediction Subtask (RQ1)
Table 5 shows the KP-PCR experimental results on the request necessity prediction compared with similarity-based and fine-tuning-based methods.
Table 5. Experiment Results of Request Necessity Prediction Subtask
Model Accuracy F1 Necessary F1 Unnecessary
CommentBERT (Ochodek et al., 2022) 0.608 0.606 0.610
MetaTransformer (Chen et al., 2022a) 0.623 0.687 0.525
Commentfinder (Hong et al., 2022) 0.616 0.711 0.426
UniPCR (Chen et al., 2024) 0.798 0.817 0.775
Our KP-PCR 0.803 0.885 0.785
In Table 5 it can be seen that KP-PCR achieves the best results compared to the other four baselines. Overall, the performance of the four methods on the non-necessary label F1 is lower than that on the necessary label F1. Based on actual data from public code reviews, this is because requests for unnecessary reviews appear in more diverse forms, such as asking questions about coding errors (which belong to coding issues rather than code review issues), submitting code that cannot be executed (which hinders reviewers from reading the code as a whole), and choosing inappropriate labels (which affects the professionalism of the review). In other words, requests that meet the review requirements have clearer expressions, executable code, and accurate identification of review issues.
(1) Compared to baseline methods, KP-PCR shows improved performance in the review necessity prediction subtask. As shown in Table 5, KP-PCR achieves a 0.7%-8.3% performance improvement over the second-best method (UniPCR) by introducing the code snippet optimization module. This demonstrates the effectiveness of our KP-PCR. By learning the representation of code snippets within the overall content through soft prompts in the prefix vectors, KP-PCR understands the differences in various code snippets under the review necessity criteria. During prediction, the representation in the learned code snippet prefix vectors enhances the prediction performance for review necessity, resulting in KP-PCR’s improved performance in this subtask.
(2) Optimizing code snippet representation is effective for the review necessity prediction task. The experimental results indicate that a large amount of representation information contained in code snippets are not fully utilized in previous traditional architectures. CommentBERT (Ochodek et al., 2022) and MetaTransformer (Chen et al., 2022a) use pre-trained language models to learn the entire request content but do not individually process or specially learn the code snippets. Commentfinder (Hong et al., 2022) treats the request body as a whole and uses similarity methods for matching. As mentioned earlier, non-essential tags in the review necessity subtask have more diverse forms and features in the data compared to essential tags, and these features are more prevalent in code snippets. The experimental results shown in Table 5 show that extracting potential features from code snippets indeed helps accomplish the review necessity prediction task.
(3) Introducing knowledge-guided prefix vectors in prompt learning is more conducive to learning code snippet representation. KP-PCR achieves an understanding of the differences in various code snippets by learning the representation of code snippets within the overall content through prefix vectors in prompt learning. This means that the small prefix vectors, during the model training process, learn the initialization parameters, allowing the relevant information for determining necessity to be extracted from the complex information contained in the overall code snippets. During prediction, the representation in the learned code snippet prefix vectors enhances the prediction performance for review necessity. Compared to the baselines, using prompt learning prefix vectors to optimize code snippets is effective.
6.2. Tag Recommendation Subtask (RQ1)
This subsection compares our KP-PCR with state-of-the-art baselines on the tag recommendation subtask and analyzes the experimental results, in terms of Precision@k, Recall@k and F1@k.
Tag recommendation subtask facilitates users to intuitively and quickly select the correct tags. As introduced in Section 3.2, users on code review websites can select 3-5 tags. Therefore, in the evaluation metrics, this study introduces @3, @5, and @10 as evaluation scopes. In addition, based on the analysis of the time complexity of our KP-PCR in Section 4.3.1, compared to the preliminary version of our study (Chen et al., 2024), our KP-PCR can improve the task accuracy performance without affecting the overall efficiency.
Table 6. Experiment Results of Tag Recommendation Subtask
Model Precision Recall F1
@3 @5 @10 @3 @5 @10 @3 @5 @10
TagDC (Li et al., 2020) 0.429 0.311 0.189 0.483 0.570 0.678 0.434 0.387 0.288
PORFIT (Nie et al., 2020) 0.511 0.368 0.217 0.571 0.668 0.774 0.516 0.456 0.330
Post2Vec (Xu et al., 2021) 0.550 0.384 0.223 0.624 0.707 0.799 0.558 0.478 0.339
PTM4Tag (He et al., 2022) 0.488 0.339 0.199 0.557 0.628 0.720 0.496 0.423 0.303
CDR4Tag (Li et al., 2023b) 0.564 0.396 0.227 0.635 0.724 0.806 0.572 0.491 0.345
UniPCR (Chen et al., 2024) 0.610 0.436 0.245 0.688 0.792 0.870 0.619 0.541 0.372
Our KP-PCR 0.613 0.438 0.247 0.692 0.793 0.871 0.622 0.543 0.373
(1) Compared to baseline methods, KP-PCR enhances the tag recommendation subtask. Our KP-PCR outperforms baselines by 0.1%-29.5% in the tag recommendation task. In terms of handling code snippets, TagDC (Li et al., 2020) and PORFIT (Nie et al., 2020) methods discard code snippets as noise, while Post2Vec (Xu et al., 2021) and PTM4Tag (He et al., 2022) methods encode code snippets separately using the same approach as text segments. The unified modeling of UniPCR adjusts the model architecture through a prompt learning paradigm. Under the similar architecture, KP-PCR achieves better performance by designing a unique training method for code snippets while maintaining the advantage of the unified modeling method, which allows for modeling multiple subtasks.
(2) Code snippet representation is also effective for the tag recommendation subtask. Unlike other baseline methods, CDR4Tag (Li et al., 2023b) designs a matching module for code snippets to capture potential information, resulting in superior performance compared to other methods. This matching module is based on similarity design, completing the task by calculating the similarity between the code snippet and the request body. However, using this similarity-based approach is still insufficient for learning code snippet representation, as the matching module can only capture parts of the code snippet that match the text content, lacking direct learning of the overall code snippet content. The experimental results indicate that focusing on learning code snippet features rather than discarding code snippet content is necessary for the tag recommendation subtask. Compared to the proposed KP-PCR, it is evident that using a training method more aligned with the structure of code snippets leads to better performance in this task.
(3) Introducing external knowledge is helpful for code snippet representation learning. Compared to UniPCR (Chen et al., 2024), it is evident that knowledge-guided prompt learning is more effective in the tag recommendation subtask. As previously mentioned, CDR4Tag (Li et al., 2023b) constructs code snippets based on a matching module, but its shortcomings lie in the complexity and difficult migration of the module, which obscures the information within the code snippets themselves. UniPCR achieves competitive experimental results by narrowing the gap through task reconstruction, while KP-PCR achieves superior performance by introducing prefix vectors and data flow graph representations of code snippets to learn the professional knowledge within the request content.
Overall, KP-PCR shows improvements across all metrics. Compared to the runner-up method UniPCR, it’s evident that knowledge-guided prompting learning is more effective in the tag recommendation subtask. Moreover, CDR4Tag constructs code snippets based on a matching module, but its drawback lies in the complexity of the module, making it difficult to transfer and obscuring the information inherent in the code snippets themselves. While KP-PCR achieves even better performance by introducing prefix vectors and data flow graph representations of code snippets for learning the professional knowledge embedded in the request content.
6.3. Effects of Knowledge-Guided Propmt Template Designs (RQ2)
This section explores experiments on the template designs for prompt learning, aiming to investigate the impact of using different designs on the model’s performance. Specifically, this section details the prompt learning designs as shown in Table 7.
(1) Knowledge-Guided Soft Prompt. Due to the abundance of domain-specific knowledge in public code reviews, this expertise is underrepresented in traditional language models due to limited coverage. Via generating domain-specific knowledge prefix vectors, external knowledge from Wikipedia is introduced to obtain professional definitions for each label. The professional definitions of the labels are extracted with a length of up to 50 words, and their representation vectors are set as the prefix vectors used in the experiments of this section.
(2) Hard Prompt Construction. Since manually constructed natural language templates are subject to human influence, for better guiding the modeling of the unified modeling method, the hard prompt template is decomposed into two parts: task definition and task description. The task definition guides the model to understand the meaning referred to by the masked word to be predicted in this task, while the task description represents the contextual background significance of the task itself. The task description indicates that the discrete prompt template only contains task-indicative text, while the task definition adds detailed explanations based on the task description. Irrelevant words are replaced with words unrelated to the task, such as “animal,” “you,” “tell,” etc., instead of the prompt language in the template.
(3) Code Prefix Prompt Construction. Code in this study is encouraged as soft prompt to learn representations through a small continuous vector, using fewer parameters to learn the overall meaning. To investigate the impact of continuous vector generation on the unified modeling method proposed in this study, this section designs experiments using irrelevant words and code snippet generation to explore the performance impact of domain-specific knowledge and prefix vector optimization. The experiment with irrelevant words uses words completely unrelated to code reviews and tasks to generate representation vectors.
Table 7. Explanation of Prompt Template Design
Name Knowledge-Guided Task Descriptive Prompt Code Prefix
(a) $\surd$ Task Description Code
(b) $\times$ Task Description Code
(c) $\times$ Irrelevant Words Code
(d) $\times$ Task Description Irrelevant Words
6.3.1. Request Necessity Prediction Subtask
Table 8 displays the experimental results on the necessity prediction subtask.
Table 8. Ablation Experiments of Request Necessity Prediction
Model Accuracy F1 Necessary F1 Unnecessary
(a) 0.830 0.885 0.785
(b) 0.798 0.817 0.775
(c) 0.745 0.771 0.712
(d) 0.775 0.782 0.767
Comparatively, KP-PCR demonstrates improved performance in the necessity prediction subtask of public code review. As shown in Table 8, compared to the suboptimal template (b), if external knowledge guidance is removed, the results will decrease by 1.3%-7.7%. Additionally, the results of template (c) and (d) compared to template (b) show that the accuracy will be reduced if the text descriptive prompt or code prefix is removed. Building upon the unified modeling proposed by KP-PCR, the soft prompt utilizes prefix vector learning to represent code snippets within the overall context, enabling an understanding of the variations in different code snippets under the review necessity criteria. During prediction, leveraging the representations from the learned code snippet prefix vectors enhances the predictive performance of review necessity, resulting in performance improvements for KP-PCR in this subtask.
6.3.2. Tag Recommendation Subtask
Table 9 displays the results of the ablation experiments on the tag recommendation subtask.
Table 9. Ablation Experiments of Tag Recommendation
Model Precision
@3 @5 @10
(a) 0.613 0.438 0.247
(b) 0.610 0.436 0.245
(c) 0.565 0.404 0.228
(d) 0.564 0.410 0.234
Model Recall
@3 @5 @10
(a) 0.692 0.793 0.871
(b) 0.688 0.792 0.870
(c) 0.671 0.772 0.853
(d) 0.628 0.736 0.826
Model F1
@3 @5 @10
(a) 0.622 0.543 0.373
(b) 0.619 0.541 0.372
(c) 0.583 0.510 0.351
(d) 0.569 0.507 0.355
(1) The introduction of external knowledge benefits the model in learning the professional knowledge contained in task data. The method of using external knowledge to initialize prefix vectors demonstrates optimal performance. As previously discussed in the analysis and description of public code review and its data in Sec 2, pre-trained language models lack knowledge of specialized terms in code review tasks (such as “c” and “psthewes”), which limits their performance in this domain. External knowledge-guided prefix tuning, by obtaining definitions of technical terms from Wikipedia and using this text as the initialization for prefix tuning, optimizes code snippet representations effectively, leading to excellent performance.
(2) It performs better by using task-related text to construct soft prompt prefix vectors. The use of unrelated words to generate prefix vectors results in the worst model performance and requires more epochs to achieve better performance in this scenario. This indicates that when initializing prefix vectors with unrelated words, the model needs more learning of its representation to adapt to the overall training objective due to a lack of guidance. Experiments initialized with code snippet representations yield suboptimal results, suggesting that initializing with task-related text enhances the model’s learning of the task. Furthermore, this also suggests that better representing code snippets can leverage more features contained in the code snippets to aid in task learning and training.
6.3.3. Discussion
(1) KP-PCR achieves better performance by optimizing the learning of code snippet representations in two quality assurance service subtasks simultaneously. By introducing external knowledge, the model can understand the specific meanings of terms in the request content. By retrieving and matching the relevant definitions from the knowledge base obtained from Wikipedia with the request content, this information is supplemented into the model to help it learn the domain-specific knowledge, resulting in a significant improvement. The experimental results demonstrate its effectiveness.
(2) External knowledge is effective for learning the semantics in the request content. Comparing the direct use of code snippet text to initialize the prefix vector part of the soft prompt, initializing the prefix vector with external knowledge leads to better performance. The use of unrelated words to generate discrete or continuous vectors has a significant impact on the overall experimental results, indicating that a prompt construction method that is more aligned with semantic understanding benefits model performance. Comparing the results of experiments using code snippet text for initialization, it can be observed that introducing external knowledge leads to performance improvement, but the overall performance enhancement is attributed to the adoption of prefix vectors as a learning method adapted to non-natural language text. The effectiveness of the unified modeling, indicating that hign performance can be achieved by selecting appropriate generation methods for initializing hard and soft prompts based on the characteristics of the data.
6.4. Case Study (RQ3)
Currently, with the progress of the GPT series models, the use of large language models to perform tasks such as summarization, translation, and dialogue has attracted widespread attention. Attributed to their powerful reasoning and generation capabilities, numerous prompt learning methods have been proposed to explore the potential of large-scale language models in practical activities, aiming to better guide them in understanding task requirements. Due to its excellent performance on various basic tasks, we design a case study to investigate the performance of our proposed prompt on large-scale language models.
First, we randomly select 20 request text and their corresponding task labels from our experimental dataset. GPT-3.5 and GPT-4.0 are then tasked to select labels based on task definitions. By altering the prompt descriptions, we investigate their performance in quality assurance tasks in public code review.
Table 10. Prompt Design for GPT-3.5 and GPT4.0
The content of the prompt template
Prompt: I need you help me to do a task which is concluding two subtasks, the one is quality prediction, and another is tag recommendation. Quality prediction is to predict the quality tag with a label “good” or “bad” for a text. The tag recommendation is to choose 10 tags in a set for each text (you can only choose tags in this set). I will give you the set for tag recommendation first, then some texts for these task.
Table 10 shows the prompt template used, while Table 11 presents the prediction results of GPT-3.5 and GPT-4.0 for the tag recommendation subtask. The bold black parts in Table 11 indicate results predicted by GPT models that are not in the real label list, while the bold blue parts represent results predicted correctly.
Table 11. The Results of Case Study
Request Text
Basic web server in Python 3. This project is intended to function as a basic web server with a command-line interface for easy use. It’s written in Python 3, with the help of the cmd module for the CLI itself and the thread module (I know, I know) to run the listening/responding process without interfering with the CLI.
I am wondering (with the exception of thread, which I do intend to fix very soon) how well it conforms to best practices and how readable it is. Performance is not a big issue.
import json
version = ””0.3.0””
settings =
serverSocket = socket.socket(socket.AF_INET, socket.SOCK_STREAM)
serverSocket.setsockopt(socket.SOL_SOCKET, socket.SO_REUSEADDR, 1)
def loadSettings():
        global settings
        with open(’.linprconfig’, ’r’) as configFile:
                settings = json.load(configFile)
Ground truth:
¡python¿¡python3x¿¡multithreading¿¡reinventing the wheel¿¡server¿
GPT-3.5:
¡python¿¡web server¿¡CLI¿¡thread module¿¡json¿¡code review¿¡socket programming¿
GPT-4.0:
¡python¿¡server¿¡python3x¿¡cpp¿¡c¿¡coding-style¿¡java¿¡unix¿¡json¿¡formatting¿
KP-PCR: ¡python¿¡beginner¿¡reinventing the wheel¿¡python3x¿¡cpp¿¡c¿¡java¿¡sqlserver¿
¡formatting¿¡server¿
It can be observed that when using natural language in defining task, GPT-3.5 and GPT-4.0 tend to generate ambiguity in the task definitions. Specifically, this manifests as returning results that do not meet the task’s requirements, such as recommending labels not in the real label list or not returning results regarding quality necessity. In contrast, when designing a more logical task statement the predicted results are more aligned with the actual task requirements.
The case study results of the tag recommendation subtask indicate that GPT demonstrates excellent performance on common knowledge labels such as programming languages (Python, Java, etc.), database knowledge (MySQL), computer systems (Linux). It can predict correctly even without pre-training. However, compared to labels containing more specialized information, GPT often fails to make accurate predictions. For example, labels such as “multithreading” and “¡reinventing the wheel” cannot be well predicted even with reference-based learning. This indicates that despite having powerful pre-trained knowledge and rapid learning abilities, large language models need to correctly understand the semantics of specialized knowledge labels.
Overall, the case study with GPT reveals that learning specialized knowledge is important for language models besides general knowledge. In the domain of public code review, there are many technical terms and information hidden within code snippets. Such characteristics make it a challenge for common large language models to handle tasks in the specialized domain of public code review.
7. Related Work
This section introduces and analyzes related work from five perspectives: tag recommendation task on software Q&A sites, public code review, prompt learning in software engineering, code language model, and comparison with our preliminary study.
7.1. Tag Recommendation Task on Software Q&A Sites
Since code reviews are conducted in the form of software Q&A sites, developers need to select appropriate tags to label their review requests during the submission stage. These tags are also used to recommend requests to suitable reviewers. Therefore, the accuracy of these tags is directly related to the final review quality.
Tag recommendation techniques have become an important research focus for software Q&A sites aiming to provide better services for code review. Numerous studies have been proposed for software information sites (Baum and Schneider, 2016). In the early stages, some research works assumed that code snippets were noise and removing them during preprocessing would improve the model’s performance such as TagDC (Li et al., 2020) and PROFIT (Nie et al., 2020).
Unlike previous works, the Post2Vec model (Xu et al., 2021) preserves code snippets and learns the content of posts, including titles, descriptions, and code snippets, seamlessly integrating them to learn a unified representation. The authors of the PTM4Tag model (He et al., 2022) argue that poorly chosen tags often introduce additional noise and redundancy, leading to problems such as tag synonyms and tag explosion. Domain knowledge is not well studied in previous studies which may reduce the quality of tag recommendation.
7.2. Public Code Review
Due to the extensive review work involved, conducting code reviews has become one of the biggest challenges facing the industry (Balachandran, 2013). Specifically, code reviews are labor-intensive because reviewers must manually inspect the code and provide written feedback. To alleviate the cost of review requests, considerable attention and research has been devoted to various automation techniques. For instance, Thongtanunam et al. (Thongtanunam et al., 2022) proposed a code transformation service that automatically converts source code into improved versions that have been reviewed and approved. Kudrjavets et al. (Kudrjavets et al., 2022) investigated opportunities for potential speedups in code by analyzing the switch from manual to automatic merging, which they found could enhance code velocity.
Unlike internal code reviews within teams, public code reviews are implemented in the form of software question-and-answer sites. This form, characterized by a large user base and rapid response, also has features such as high knowledge dissemination. However, it also introduces factors that can affect quality and efficiency, primarily related to developers. Therefore, with the popularity of public code review, it has led to task demands similar to code review to ensure the quality of reviews.
As shown in Table 12, previous relevant studies have relied mainly on internal team data or other thematic site data to accomplish intelligent and automated construction of these two tasks. Since the quality and responsiveness of a public code review are typically determined by developers, this process is often carried out manually, making it difficult to achieve convenience and quality assurance. Therefore, with the popularity of this review format, the demand for a shift from manual to intelligent services in public code reviews has also arisen. The process of ensuring review quality in public code reviews can be seen as a review necessity prediction and a tag recommendation task.
Table 12. Popular Software Engineering Thematic Sites
Name Thematic
StackOverFlow Programming Q&A Exchange Sites
WordPress Open-source Blog Design Sites
Code Review Public Code Review Sites
Github Software Project Hosting Sites
Database Administrator Database Management Q&A Sites
Li et al. (Li et al., 2022) proposed that by systematically considering the tasks involved in the entire review process and modeling them, the time and resource costs of the review process can be greatly reduced. Existing research works mostly employ fine-tuning paradigm modifications of model architectures to ensure the model’s performance in terms of accuracy and other relevant metrics for the tasks. This paradigmatic method can accomplish individual tasks with individual models. Table 13 shows the statistical information on model construction for the aforementioned two subtasks. It can be observed that although different modeling methods are used for different tasks, most existing research is based on pre-trained language models. Therefore, considering that existing works in both subtasks have mostly used pre-trained language models for construction in recent years, it is worth considering designing a unified modeling method to simultaneously address both subtasks and reduce cost requirements.
Table 13. The Model Classifications Used in Existing Methods
Name Pre-trained Language CNN DAG Similarity
TagDC (Li et al., 2020) $\surd$
PORFIT (Nie et al., 2020) $\surd$
CommentBERT (Ochodek et al., 2022) $\surd$
MetaTransformer (Chen et al., 2022a) $\surd$
Commentfinder (Hong et al., 2022) $\surd$
Post2Vec (Xu et al., 2021) $\surd$
PTM4Tag (He et al., 2022) $\surd$
CDR4Tag (Li et al., 2023b) $\surd$
UniPCR (Chen et al., 2024) $\surd$
7.3. Prompt Learning in Software Engineering
Fine-tuning paradigm, where models tailored to specific tasks are trained only on target task’s input-output example datasets, has long played a central role in many machine learning tasks (Lu et al., 2023; Han et al., 2022; Li and Liang, 2021). Currently, prompt learning can be divided into two types: hard prompts and soft prompts.
Currently, both have found rich applications in various natural language processing tasks. For example, Han et al. (Han et al., 2022) augmented MLM with rule-based designs to decompose challenging tasks into several simpler sub-tasks. Similarly, Li et al. (Li and Liang, 2021) introduced prefix tuning, which optimizes small vectors of task-specific parameters, achieving similar performance to fine-tuning by learning only 0.1% of the parameters.
In the field of software engineering, prompt tuning has seen increasing applications. Due to the lightweight and efficient nature pursued in software engineering tools, existing research efforts have gradually attempted to use methods similar to prompt tuning as alternatives to fine-tuning. For example, Lu et al. (Lu et al., 2023) proposed an LLaMA-Reverwer framework, which leverages mainstream large language models and adopts a parameter-efficient fine-tuning method to address the computational challenges of fine-tuning large language models. Similarly, the latest tool, POME, introduced by Huang et al. (Huang et al., 2023), infers non-fully qualified type name syntax and usage knowledge encapsulated in prompt-tuned code-masked language models through a fill-in-the-blank strategy.
7.4. Code Language Model
To better capture the inherent structure of code, GraphCodeBERT proposed by Guo et al. (Guo et al., 2021) and TRACED proposed by Ding et al. (Ding et al., 2024) can be used. GraphCodeBERT demonstrated significant improvements in adapting to downstream tasks compared to other pre-trained language models across multiple code-related tasks. GraphCodeBERT, as such a programming language-based pre-trained model, is constructed inspired by natural language pre-trained models, enhancing model training by incorporating code-specific pre-training tasks on top of natural language pre-training. Existing pre-trained models treat code snippets as token sequences, ignoring the inherent structure in code that provides crucial semantics for enhanced code understanding. TRACED, an execution-aware pre-training strategy for source code, to teach code models the complicated execution logic during the pre-training, enabling the model to statically estimate the dynamic code properties without repeatedly executing code during task-specific fine-tuning. GraphCodeBERT utilizes data flows in the pre-training phase, representing the semantic-level structure of code, encoding relationships between variable values. This semantic-level structure is less complex and does not introduce unnecessary deep hierarchical structures, making the model more effective. GraphCodeBERT is used as pre-trained model in our study.
7.5. Comparison with Our Preliminary Study
In our preliminary study (Chen et al., 2024), a unified model is proposed for the request necessity prediction subtask and the tag recommendation subtask. The study in this paper represents a significant expansion of that study in several respects. Firstly, we introduce an external knowledge base by soft prompt to generate prefix vectors which guide the model to understand the meaning of the code snippet. Secondly, we design the knowledge-guided prompt learning on request quality assurance task to show that the knowledge guide can further improve the request quality of PCR. Thirdly, we set up a prompt template design study to explore the influence of external knowledge. Further, we analyze the time complexity to prove that our KP-PCR can improve task accuracy while maintaining overall efficiency. Finally, we design a case study to investigate the performance of our prompt templates on large-scale language models. In accordance with these additional results, the entire paper has been thoroughly updated to reflect the new holistic study.
8. Conclusions and Future Work
Through investigating and summarizing the relevant tasks of team code review and software question-and-answer sites, and combining them with the process of public code review, this study explores intelligent service methods for the necessity prediction task and tag recommendation task of public code review from the perspective of developers. Addressing the issue of semantic understanding of professional knowledge in the request content, we introduce external knowledge and adopt a processing and training method that conforms to the semantic structure of code snippets to achieve a more comprehensive learning of professional knowledge. We propose KP-PCR to optimize the learning of professional knowledge in public code review. In public code review, request content includes both text and code snippets. The text contains vocabulary related to professional knowledge, while the unique structure of code snippets contains representations of content strongly related to review quality as the main part of the content. By processing code snippets into data flow graphs and guiding the model to learn professional knowledge with knowledge-guided prefix vectors, KP-PCR achieves significant improvements in both subtasks. In the experimental results of the two subtasks, KP-PCR shows an improvement of 8.3% to 28.8% in the necessity prediction subtask and an improvement of 0.1% to 29.5% in the tag recommendation subtask.
This study has used Wikipedia as a source of external knowledge, and the influence of more different knowledge repositories can be explored in the future. In this study, the model was implemented based on the pre-trained language model BERT, which performed well among baseline methods. Further, our prompt learning can be transferred to more large language models in the future.
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant Nos. 62276196 and 62172311.